Shomika was helping her family pick oranges in their grove. She took some oranges home to share with her three friends. She gave 3 more than half to Jennifer. Angela got half of the remainder and 3 more. She gave Josie half of the remainder plus 3. When she got home, she only had 10 oranges left. How many did she have when she left the grove?
Answer by stanbon(75887)

Shomika was helping her family pick oranges in their grove.
She took some oranges home to share with her three friends.
-----------
Let number she started with be "x".
------------------------
She gave 3 more than half to Jennifer.
She now has (1/2)x-3 = [(x-6)/2] oranges:
---------------------
Angela got half of the remainder and 3 more.
Now has (1/2)[(x-6)/2] -3 = (x-6)/4 - 12/4 = [(x-18)/4]
-----------------------------
She gave Josie half of the remainder plus 3.
Now has (1/2)[x-18)/4] -3 = [(x-18)/8 - 24/8] = (x-42)/8
----------------------------------
When she got home,she only had 10 oranges left.
How many did she have when she left the grove?
EQUATION:
(x-42)/8 = 10
x-42 = 80
x = 122 oranges (the numbers she started with.
=========================
Cheers,
Stan H.